# Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior

### v4

@@ V Impact of Customers Strategic Behavior
In this section, we analyze gaming agents’ strategic behavior in the two-agent two-period setting. First, we state a benchmark centralized CP shaving model for comparison, then analyze agents’ strategic behavior from both an economic perspective with the PoA, and a technical perspective with the peak shaving effectiveness.
@@ V-A Centralized CP shaving
We first state the centralized CP shaving model, which assumes a central operator has direct control over both agents in the formulated two-period condition to reduce the total cost of both agents. This centralized model thus maximizes the total objective function of both agents and is denoted as
@@
We now present the following proposition, demonstrating that the centralized CP shaving model is equivalent to the centralized peak shaving model. This model minimizes the total peak demand of both agents while also incorporating peak demand charges and shifting costs.
@@
Centralized peak shaving model. The centralized CP shaving model (14) is equivalent to the trivial convex peak shaving model as follows:
$(x^{*},y^{*})\in\arg\min_{x,y}\ \pi\max\{S_{1}(x,y),S_{2}(x,y)\}+\alpha_{\mathrm{x}}x^{2}+\alpha_{\mathrm{y}}y^{2}$ (15)
@@
$f_{\mathrm{x}}(x,y)$ $+f_{\mathrm{y}}(x,y)=\pi(X_{1}+x)I(S_{1}(x,y)-S_{2}(x,y))+\pi(X_{2}-x)I(S_{2}(x,y)-S_{1}(x,y))$
$+\pi(Y_{1}+y)I(S_{1}(x,y)-S_{2}(x,y))+\pi(Y_{2}-y)I(S_{2}(x,y)-S_{1}(x,y))+\alpha_{\mathrm{y}}y^{2}+\alpha_{\mathrm{x}}x^{2}$
$=\pi S_{1}(x,y)I(S_{1}(x,y)-S_{2}(x,y))+\pi S_{2}(x,y)I(S_{2}(x,y)-S_{1}(x,y))+\alpha_{\mathrm{y}}y^{2}+\alpha_{\mathrm{x}}x^{2}$
$=\pi\max\{S_{1}(x,y),S_{2}(x,y)\}+\alpha_{\mathrm{y}}y^{2}+\alpha_{\mathrm{x}}x^{2}.$ (16)
Note that the peak shaving model in (15) is convex because maximize two convex (linear) functions $S_{1}(x,y),S_{2}(x,y)$ is convex; thus, we prove the Proposition. ∎
This model assumes direct control over each agent’s demand, achieving peak shaving at minimal cost. However, while its performance is strong, the assumption of direct load control is unrealistic in practice, as agents (customers) typically have independent strategies and privacy concerns. Therefore, we consider (15) as a benchmark to compare with our CP game model and analyze peak shaving effectiveness and the PoA in the following sections.
@@ V-B Peak shaving ratio analysis
We show in the following theorem that the CP game can achieve the same effectiveness in reducing the system peak demand at its equilibrium.
@@
Peak shaving performance. The peak shaving effectiveness of the CP game model $G$ at equilibrium is always 1, i.e.,
@@
We prove this Theorem by first analytically writing the solution for the centralized model shown in Proposition 8, and from Theorem 5, we have the game model solution. Then take the solutions of the centralized model and the game model under the same condition into the peak shaving effectiveness definition in this Theorem. The detailed proof is provided in the appendix. ∎
This Theorem shows the game model always reaches the same peak shaving performance when compared with the centralized model. The reason is that agents in both the game model and centralized model shift demand as much as possible to avoid CP charge without burdening more by their shifting penalty. Specifically, under concave game conditions, both agents’ shifting capability is constrained by their critical points, preventing them from fully balancing their demand—similar to the centralized model. Under quasiconcave and non-concave game conditions, system demand is balanced across the two periods, also aligning with the centralized model. This suggests that the game model is practically applicable, as utilities and operators can implement the game framework for CP shaving without compromising peak shaving performance. In other words, operators can still achieve their peak shaving targets while using the game model. However, although the overall peak shaving performance is the same, agents’ individual demand shifting is different due to the information barrier, which reflects as cost of reaching the peak shaving performance. We then analyze the cost by showing the PoA in the following section.
V-C PoA analysis
In this section, we study the PoA affected by gaming agents’ strategic behavior in all three game structures, and we provide the following main results.
@@ V-C PoA analysis
PoA with agent equity. Given the PoA defined as
@@
under the quasiconcave and non-concave game condition as described in Proposition 4, the PoA increases with the disparities among agents, as measured by the marginal shifting cost $\alpha_{\mathrm{x}}x^{*},\alpha_{\mathrm{y}}y^{*}$, i.e.,
@@
From Theorem 5, we know the NE of the game model under quasiconcave and non-concave game conditions. Combined with the centralized model solution obtained from Theorem 9, we show the difference of nominator and denominator of the PoA as defined in (18a) can be expressed as a Euclidean distance between agents’ marginal shifting cost, i.e., $(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}$. The detail is provided in the appendix. ∎
Under non-concave and quasiconcave game conditions, although the overall CP charge is always the same as the centralized model due to the same peak shaving effectiveness, the shifting cost $\alpha_{\mathrm{x}}x^{*2}+\alpha_{\mathrm{y}}y^{*2}$ increases due to the information barrier. We show in this theorem that the PoA increases with agents’ disparities, quantified by their marginal shifting cost. This insight provides a pathway for designing mechanisms that balance agents’ marginal shifting costs to ensure both system effectiveness and fairness. For instance, agents with both a large shifting penalty parameter and a high shifting amount bear a higher marginal shifting cost. Here, the penalty parameter reflects greater comfort loss from demand shifting, while the high shifting amount arises from significant demand differences requiring adjustment due to CP charges. In such cases, operators or utilities could regulate these agents to shift less demand or provide incentives to reduce their penalty parameters. As some policies already suggest subsidizing disadvantaged customers, this analysis offers guidance on how to allocate subsidies effectively among customers [Pacific Gas & Electric Company, 2025, Consolidated Edison, 2025b]. By balancing marginal shifting costs, this approach simultaneously improves system equity and reduces the PoA.
@@
PoA with CP game type. For given $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$, the PoA defined as Theorem 10 is highest at equilibrium for quasiconcave games, followed by non-concave games, and lowest for concave games, where it is always equal to 1, i.e.,
@@
We first show $P=1$ is always true for the concave game due to the same solution structure between the centralized model and concave game model based on the results from Theorems 5 and 9. Then from Theorem 10, we know the PoA expression as Euclidean distance between agents’ marginal shifting cost under quasiconcave and non-concave game conditions, indicating $P\geq 1$. Thus, quasiconcave games and non-concave games always cause higher (or equal) PoA than concave games. We then show by fixing $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$, the agent’s baseline demand $X_{1},X_{2},Y_{1},Y_{2}$ can vary to cause different game structures, so as to different solution structure $x^{*},y^{*}$ from Theorems 5. Combined with the fact that quasiconcave and non-concave games always balance system demand, i.e., $x^{*}+y^{*}=b$ from Theorem 5, we prove the quasiconcave game solution shows a higher difference between $x^{*},y^{*}$, indicating higher PoA than a non-concave game. The detailed proof is provided in the appendix. ∎
@@
This theorem demonstrates the impact of the CP game type on the PoA, where the game type connects to the agents’ flexibility. We first fix the system conditions $\pi,S$ while allowing the agents’ conditions to vary. From Theorem 5, we know that agents’ flexibility is determined by their shifting penalty parameters $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$ and baseline demand $X_{1},X_{2},Y_{1},Y_{2}$. To control the influence of both parameters, we also fix $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$, making the PoA dependent only on $x^{*},y^{*}$, which reflects the difference in the game structure. As the balance point (demand difference) $b_{\mathrm{x}}=(X_{2}-X_{1})/2$ increases, agent x’s flexibility decreases, and the game type transitions from a quasiconcave game to a non-concave game, and finally to a concave game. This follows from the conditions $\{-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}\}\cap\{-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}}\}$ to $\{-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}\}\cap\{b\geq r_{\mathrm{x}}+r_{\mathrm{y}}\cup b\leq-r_{\mathrm{x}}-r_{\mathrm{y}}\}$ to $\{r_{\mathrm{x}}+r_{\mathrm{y}}<b\}\cup\{-r_{\mathrm{x}}-r_{\mathrm{y}}>b\}$ as stated in Proposition 4. Thus, the PoA increases with agents’ flexibility and is also reflected in the game type change. This highlights the role of inflexible agents in overall system performance, despite utilities and operators generally favoring flexible agents for their ability to enhance grid stability, integrate renewable energy, and respond to pricing signals [Siano, 2014]. It is also evident that fixing $X_{1},X_{2},Y_{1},Y_{2}$ while allowing $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$ to vary yields the same results. Note that this theorem also shows that under the concave game condition, the system’s performance is always equivalent to that of the centralized model, indicating no harm to utility companies and agents when applying the concave CP game.
@@ VI Extension to Multi-agent Games
In this section, we consider a generalization of our model from the two-agent two-period CP game to the multi-agent two-period CP game, and we analyze the NE, stability/convergence, as well as the consumers’ strategic behavior.
VI-A Multi-agent CP game
We denote $X_{i,1},X_{i,2}$ as agent $i$’s baseline demand in period 1 and 2, $\alpha_{i}$ as agent $i$’s penalty parameter, and $x_{i}$ as agent $i$’s demand shifting strategy. We then extend the CP game $G$ from two-agent to multi-agent with $G^{\prime}=(N,\mathcal{X},U)$, such that,
  • $N=\{1,2,...,|N|\}$ with index $i$ is the agent set, where ||$\cdot$|| ⋅ | for a set means the number of elements inside a set (otherwise, it is absolute value); specifically, we use $-i$ to indicate all agents but $i$.
@@ VI-A Multi-agent CP game
$\max_{x_{i}}$ $f_{i}(x_{i},x_{-i})=-\pi(X_{i,1}+x_{i})I(S_{1}(x_{i},x_{-i})-S_{2}(x_{i},x_{-i}))$
$-\pi$ $(X_{i,2}-x_{i})I(S_{2}(x_{i},x_{-i})-S_{1}(x_{i},x_{-i}))-\alpha_{i}x_{i}^{2},$ (19a)
$S_{1}(x_{i},x_{-i})$ $=\sum_{i\in N}(X_{i,1}+x_{i})=S_{\mathrm{b},1}+\sum_{i\in N}x_{i},$ (19b)
$S_{2}(x_{i},x_{-i})$ $=\sum_{i\in N}(X_{i,2}-x_{i})=S_{\mathrm{b},2}-\sum_{i\in N}x_{i},$ (19c)
Note that the system demands $S_{1},S_{2}$ change from the summation of two agents’ to $N$ agents’ demand at periods 1 and 2.
The multi-agent CP game $G^{\prime}$ still follow the property as described in Proposition 4, i.e., the CP game $G^{\prime}$ is
@@
    $b<-\sum_{i\in N}r_{i},\text{when }S_{\mathrm{b},1}\geq S_{\mathrm{b},2},b>\sum_{i\in N}r_{i},\text{when }S_{\mathrm{b},1}<S_{\mathrm{b},2},$ (20a)
@@
    $\{-\sum_{i\in N}r_{i}\leq b\leq\sum_{i\in N}r_{i}\}\cap\{\exists i\in N,b_{i}<-r_{i}\cup b_{i}>r_{i}\}.$ (20c)
The concave and quasiconcave conditions are intuitive and can be obtained following the same process of Proposition 4. Then the non-concave condition is determined by the complementary set of concave and quasiconcave conditions.
In terms of the NE solutions, we can’t analyze each agent’s capability case by case like the two-agent setting. Thus, we separately analyze NE in each game type. We first show concave game has a unique NE according to [Rosen, 1965], where the NE definition slightly changes from Definition 1 to include all agents, i.e.,
@@
The NE is then obtained when all agents reach their critical point as defined in Definition 2 corresponding to the active periods, and we have the NE as follows,
$x_{i}^{*}$ $=-r_{i},i\in N,\text{if }b<-\sum_{i\in N}r_{i},$ (22a)
$x_{i}^{*}$ $=r_{i},i\in N,\text{if }b>\sum_{i\in N}r_{i}.$ (22b)
Noted that when all $x_{i}^{*}$ take $-r_{i}$, the CP period is always period 1 and $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$, and agents shift demand from period 1 to period 2, until reaching their critical point.
NE of the quasiconcave game is not intuitive and we introduce the following Proposition to analyze it.
@@
Existence and uniqueness of NE in multi-agent quasiconcave CP game. The quasiconcave CP game $G^{\prime}$ as described in (19) and satisfy (20b) has a unique pure-strategy NE $(x_{i}^{*},x_{-i}^{*})$ as defined in (21), where
@@
To prove this Proposition, we derive two Lemmas showing the existence (Lemma 17) and uniqueness (Lemma 18). The basic principle of existence Lemma is based on [Reny, 1999], i.e., the sum of the agent’s payoff functions is upper semi-continuous in $x_{i}\in\mathcal{X}$ and game $G^{\prime}$ is payoff security, where we define in the proof. The proof of uniqueness Lemma is based on [Rosen, 1965]. We first prove the uniqueness of NE in periods 1 and 2. Then, we start with a two-agent setting to analyze all agents’ best responses and use a min-max formulation to show the NE of each agent. After that, we apply these two-agent results to multi-agent settings by sequentially partitioning agents into two groups. The detailed proof is provided in the appendix. ∎
This Proposition shows the quasiconcave CP game $G^{\prime}$ has a unique NE and always obtains at their balance points $b_{i}$. Note that under both concave and quasiconcave conditions, the NE results remain consistent with the two-agent setting, as the extension from two agents to multiple agents follows a straightforward generalization. This suggests an approach where the NE results of the two-agent non-concave CP game $G$ serve as a foundation for analyzing the $N$-agent non-concave CP game $G^{\prime}$, whose NE is evidently non-trivial. Basically, Theorem 5 conveys an important message that the system demand will always be balanced in the two periods, i.e., $x^{*}+y^{*}=b$, and one agent (less flexible according to shifting penalty parameters and baseline demand) first reach the critical point while the other agent balance the system demand in the two periods.
Here we denote the agents whose baseline peak demand is in the system baseline CP period as CP-period agent, and the other as non-CP-period agent. We also call the set that includes all CP-period agents as CP-period agent set and includes all non-CP-period agents as non-CP-period agent set, denote as $N_{\mathrm{cp}},N_{\mathrm{ncp}}$, respectively, such that $N_{\mathrm{cp}}\cup N_{\mathrm{ncp}}=N,N_{\mathrm{cp}}\cap N_{\mathrm{ncp}}=\emptyset$. We then provide the following Proposition to analyze the equilibrium condition of the non-concave multi-agent CP game $G^{\prime}$.
@@
NE in non-concave multi-agent CP game. Given non-concave multi-agent CP game $G^{\prime}$ and $S_{\mathrm{b},1}<S_{\mathrm{b},2}$, under the condition of
@@
the hybrid NE solution, i.e., for each CP-period agent and for the whole non-CP-period agent set, is given by
$x_{i}^{*}=\min\{r_{i},b_{i}\},i\in N_{\mathrm{cp}},\sum_{i\in N_{\mathrm{ncp}}}x_{i}^{*}=b-\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\};$ (24b)
otherwise, the hybrid NE solution, i.e., for each non-CP-period agent and the overall CP-period agent set, is given by
@@
The basic idea is to assume two virtual agents that have the same baseline demand and shifting penalty conditions with the CP-period agent set and non-CP-period agent set. However, their strategy structures are different, where the agent set’s strategy is determined by the aggregation of each individual agent’s strategy. From Theorem 5, we derive the best strategies of two virtual agents, and from the best response analysis, we know the agent sets and virtual agents have the same response rationale due to the same baseline conditions. Basically, they will both shift demand away from the CP period or to the CP period to avoid the cost, and the cost reduction positively relates to their shifting amount. We then link the two virtual agents’ strategies to the agent sets’ strategy and show the equilibrium point for the CP-period agent set and the non-CP-period agent set under different baseline conditions. The detailed proof is provided in the appendix. ∎

### v5

@@ V Impact of Customers Strategic Behavior
In this section, we analyze gaming agents’ strategic behavior in the two-agent two-period setting. First, we state a benchmark centralized CP shaving model for comparison, then analyze agents’ strategic behavior from both an economic perspective with the efficiency loss, and a technical perspective with the peak shaving effectiveness.
@@ V-A Centralized CP shaving
We first introduce the centralized CP shaving model, which assumes a central operator has direct control over both agents in the formulated two-period setting to minimize their total cost. This model represents a centralized realization of the CP shaving game. Consequently, the centralized model maximizes the total objective function of both agents and is formulated as follows
@@
We now present the following proposition, demonstrating that the centralized CP shaving model is equivalent to a travail centralized peak shaving model. This model minimizes the total peak demand of both agents while also incorporating peak demand charges and shifting costs.
@@
Centralized peak shaving model. The centralized CP shaving model (14) is equivalent to the trivial concave peak shaving model as follows:
$(x^{*},y^{*})\in\arg\max_{x,y}\ -\pi\max\{S_{1}(x,y),S_{2}(x,y)\}-\alpha_{\mathrm{x}}x^{2}-\alpha_{\mathrm{y}}y^{2}$ (15)
@@
$f_{\mathrm{x}}(x,y)$ $+f_{\mathrm{y}}(x,y)=-\pi(X_{1}+x)I(x,y)-\pi(X_{2}-x)(1-I(x,y))$
$-\pi(Y_{1}+y)I(x,y)-\pi(Y_{2}-y)(1-I(x,y))-\alpha_{\mathrm{y}}y^{2}-\alpha_{\mathrm{x}}x^{2}$
$=-\pi S_{1}(x,y)I(x,y)-\pi S_{2}(x,y)(1-I(x,y))-\alpha_{\mathrm{y}}y^{2}-\alpha_{\mathrm{x}}x^{2}$
$=-\pi\max\{S_{1}(x,y),S_{2}(x,y)\}-\alpha_{\mathrm{y}}y^{2}-\alpha_{\mathrm{x}}x^{2}.$ (16)
Note that the peak shaving model in (15) is concave because maximize two concave (linear) functions $S_{1}(x,y),S_{2}(x,y)$ is concave; thus, we prove the Proposition. ∎
This model assumes direct control over each agent’s demand, achieving peak shaving at minimal cost. However, while its performance is strong, the assumption of direct control over each agents’ demand is unrealistic in practice, as agents (customers) typically have independent strategies and privacy concerns. Therefore, we consider (15) as a benchmark to compare with our CP shaving game model and analyze peak shaving effectiveness and efficiency loss in the following sections.
@@ V-B Peak shaving ratio analysis
We show in the following theorem that the CP shaving game can achieve the same effectiveness in reducing the system peak demand at its equilibrium.
@@
Peak shaving performance. The peak shaving effectiveness of the CP shaving game model $G$ at equilibrium is always 1, i.e.,
@@
We prove this Theorem by first analytically writing the best solution for the centralized model shown in Proposition 2, and from Theorem 1, we have the game model equilibrium. Then take the solutions of the centralized model and the game model under the same condition into the peak shaving effectiveness definition in this Theorem. The detailed proof is provided in the appendix. ∎
This Theorem shows the game model always reaches the same peak shaving performance when compared with the centralized model. The reason is that agents in both the game model and centralized model shift demand as much as possible to avoid CP charge without burdening more by their shifting penalty. Specifically, under concave game conditions, both agents’ shifting capability is constrained by their critical points, preventing them from fully balancing their demand—similar to the centralized model. Under quasiconcave and non-concave game conditions, system demand is balanced across the two periods, also aligning with the centralized model. This suggests that the game model is practically applicable, as utilities and operators can implement the game framework for CP shaving without compromising peak shaving performance. In other words, operators can still achieve their peak shaving targets while using the game model. However, although the overall peak shaving performance is the same, agents’ individual demand shifting is different due to the information barrier, which reflects as cost of reaching the peak shaving performance. We then analyze the cost by showing the efficiency loss in the following section.
V-C Efficiency loss analysis
In this section, we study the efficiency loss affected by gaming agents’ strategic behavior in all three game structures, and we provide the following main results.
@@ V-C Efficiency loss analysis
Efficiency loss with agent equity. Given the efficiency loss defined as
@@
under the quasiconcave and non-concave game condition as described in Proposition 1, the efficiency loss increases with the disparities among agents, as measured by the marginal shifting cost $\alpha_{\mathrm{x}}x^{*},\alpha_{\mathrm{y}}y^{*}$, i.e.,
@@
From Theorem 1, we know the NE of the game model under quasiconcave and non-concave game conditions. Combined with the centralized model solution obtained from Theorem 4, we show the difference of nominator and denominator of the efficiency loss as defined in (18a) can be expressed as a Euclidean distance between agents’ marginal shifting cost, i.e., $(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}$. The detail is provided in the appendix. ∎
Under non-concave and quasiconcave game conditions, although the overall CP charge is always the same as the centralized model due to the same peak shaving effectiveness, the shifting cost $\alpha_{\mathrm{x}}x^{*2}+\alpha_{\mathrm{y}}y^{*2}$ increases due to the information barrier. We show in this theorem that the efficiency loss increases with agents’ disparities, quantified by their marginal shifting cost. This insight provides a pathway for designing mechanisms that balance agents’ marginal shifting costs, which also enhances fairness while improving overall system effectiveness. For instance, agents with both a large shifting penalty parameter and a high shifting amount bear a higher marginal shifting cost. Here, the penalty parameter reflects greater comfort loss from demand shifting, while the high shifting amount arises from significant demand differences requiring adjustment due to CP charges. In such cases, operators or utilities could regulate these agents to shift less demand or provide incentives to reduce their penalty parameters. As some policies already suggest subsidizing disadvantaged customers, this analysis offers guidance on how to allocate subsidies effectively among customers [Pacific Gas & Electric Company,, 2025; Consolidated Edison, 2025b, ]. By balancing marginal shifting costs, this approach simultaneously improves system fainness and efficiency.
@@
Efficiency loss with CP shaving game type. For given $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$, the efficiency loss defined as Theorem 5 is highest at equilibrium for quasiconcave games, followed by non-concave games, and lowest for concave games, where it is always equal to 1, i.e.,
@@
We first show $P=1$ is always true for the concave game due to the same solution structure between the centralized model and concave game model based on the results from Theorems 1 and 4. Then from Theorem 5, we know the efficiency loss expression as Euclidean distance between agents’ marginal shifting cost under quasiconcave and non-concave game conditions, indicating $P\geq 1$. Thus, quasiconcave games and non-concave games always cause higher (or equal) efficiency loss than concave games. We then show by fixing $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$, the agent’s baseline demand $X_{1},X_{2},Y_{1},Y_{2}$ can vary to cause different game structures, so as to different solution structure $x^{*},y^{*}$. Combined with the fact that quasiconcave and non-concave games always balance system demand, i.e., $x^{*}+y^{*}=b$ from Theorem 1, we prove the quasiconcave game solution shows a higher difference between $x^{*},y^{*}$, indicating higher efficiency loss than a non-concave game. The detailed proof is provided in the appendix. ∎
@@
This theorem demonstrates the impact of the CP shaving game type on the efficiency loss, where the game type connects to the agents’ flexibility (shifting capability). We first fix the system conditions $\pi,S$ while allowing the agents’ conditions to vary. From Theorem 1, we know that agents’ flexibility is shaped by the game structure, which in turn is influenced by the shifting penalty parameters $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$ and baseline demand conditions $X_{1},X_{2},Y_{1},Y_{2}$. To control the influence these parameters, we also fix $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$, making the efficiency loss dependent only on $x^{*},y^{*}$, which reflects the difference in the game structure. As the balance point (demand difference) $b_{\mathrm{x}}=(X_{2}-X_{1})/2$ increases, agent x’s flexibility decreases, and the game type transitions from a quasiconcave game to a non-concave game, and finally to a concave game. This follows from the conditions $\{0\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}\}\cap\{-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}}\}$ to $\{0\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}\}\cap\{b_{\mathrm{x}}>r_{\mathrm{x}}\cup b_{\mathrm{y}}<-r_{\mathrm{y}}\cup b_{\mathrm{y}}>r_{\mathrm{y}}\}$ to $\{r_{\mathrm{x}}+r_{\mathrm{y}}<b\}$ as stated in Proposition 1. Thus, the efficiency loss increases with agents’ flexibility and is also reflected in the game type change. This highlights the role of inflexible agents in overall system performance, despite utilities and operators generally favoring flexible agents for their ability to enhance grid stability, integrate renewable energy, and respond to pricing signals [Siano,, 2014].
It is also evident that fixing $X_{1},X_{2},Y_{1},Y_{2}$ while allowing $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$ to vary yields the same results. Note that this theorem also shows that under the concave game condition, the system’s performance is always equivalent to that of the centralized model, indicating no harm to utility companies and agents when applying the concave CP shaving game.
@@ VI Extension to Multi-agent Games
In this section, we consider a generalization of our model from the two-agent two-period CP shaving game to the multi-agent two-period CP shaving game, and we analyze the NE, stability/convergence, as well as the customers’ strategic behavior.
VI-A Multi-agent CP shaving game
We denote $X_{i,1},X_{i,2}$ as agent $i$’s baseline demand in periods 1 and 2, respectively, and extend the system baseline demand $S_{\mathrm{b},1},S_{\mathrm{b},2}$ from the sum of two agents’ demands to the sum over $N$ agents at each period. Without loss of generality, We maintain the assumption that the baseline demand in period 2 is greater than in period 1, i.e., $S_{\mathrm{b},1}<S_{\mathrm{b},2}$. We use $\alpha_{i}$ to denote agent $i$’s penalty parameter, and $x_{i}$ as its demand-shifting strategy. Note that we do not impose any assumptions on individual agents’ baseline demands, ensuring that all $x_{i}$ remain unrestricted. We continue to use $S_{1},S_{2}$ to represent the total system demand, incorporating all agents’ demand shifts. The multi-agent game model is then formally defined as $G^{\prime}=(N,\mathcal{X},U)$, where
  • $N=\{1,2,...,|N|\}$ with index $i$ is the agent set, where ||$\cdot$|| ⋅ | represents the number of elements in a set (otherwise, it denotes the absolute value). Specifically, we use $-i$ to refer to all agents except agent $i$.
@@ VI-A Multi-agent CP shaving game
$\max_{x_{i}}f_{i}(x_{i},x_{-i})$ $=-\pi(X_{i,1}+x_{i})I(x_{i},x_{-i})-\pi(X_{i,2}-x_{i})(1-I(x_{i},x_{-i}))-\alpha_{i}x_{i}^{2},$ (19a)
$I(x_{i},x_{-i})$ =$=\begin{cases}1&S_{1}(x_{i},x_{-i})-S_{2}(x_{i},x_{-i})\geq 0\\ 0&S_{1}(x_{i},x_{-i})-S_{2}(x_{i},x_{-i})<0\end{cases},$ (19b)
$S_{1}(x_{i},x_{-i})$ $=\sum_{i\in N}(X_{i,1}+x_{i})=S_{\mathrm{b},1}+\sum_{i\in N}x_{i},$ (19c)
$S_{2}(x_{i},x_{-i})$ $=\sum_{i\in N}(X_{i,2}-x_{i})=S_{\mathrm{b},2}-\sum_{i\in N}x_{i},$ (19d)
Accordingly, the system average demand $S$ and the balance point $b$ as defined in Definition 2 are extended from the two-agent case to the $N$ agent case. Thus, we can conclude that the multi-agent CP shaving game $G^{\prime}$ retains the properties described in Proposition 1, meaning that the CP shaving game $G^{\prime}$ is
@@
    $b=\sum_{i\in N}b_{i}>\sum_{i\in N}r_{i}>0;$ (20a)
@@
    $\{0\leq b\leq\sum_{i\in N}r_{i}\}\cap\{\exists i\in N,b_{i}<-r_{i}\cup\exists i\in N,b_{i}>r_{i}\}.$ (20c)
The concave and quasiconcave conditions are intuitive and can be derived following the same process outlined in Proposition 1. The non-concave condition, on the other hand, is determined by the complementary set of the concave and quasiconcave conditions with respect to $\mathbb{R}$.
In terms of NE solutions, the definition of NE should be slightly modified from Definition 1 to account for all agents. Specifically, $(x^{*},x_{-i}^{*})$ is a pure strategy NE of the game $G^{\prime}$ if and only if
@@
Unlike the two-agent setting, analyzing each agent’s capability case by case becomes infeasible in the multi-agent setting. Therefore, we analyze the NE separately for each game type. First, for the concave game, we establish the existence of a unique NE following [Rosen,, 1965]. The NE is reached when all agents shift demand up to their critical points, as defined in Definition 2:
$x_{i}^{*}$ $=r_{i},i\in N.$ (22)
Noted that when all agents adopt $x_{i}^{*}=r_{i}$, the CP period remains period 2 and $S_{\mathrm{b},1}<S_{\mathrm{b},2}$. In this case, all agents shift demand from period 2 to period 1 until reaching their respective critical points.
The NE of the quasiconcave game is less straightforward to derive. To analyze it rigorously, we introduce the following proposition.
@@
Existence and uniqueness of NE in multi-agent quasiconcave CP shaving game. The quasiconcave CP shaving game $G^{\prime}$ as described in (19) and satisfy (20b) has a unique pure-strategy NE $(x_{i}^{*},x_{-i}^{*})$ as defined in (21), where
@@
To prove this proposition, we establish two supporting lemmas: one for the existence of NE (Lemma 3) and another for its uniqueness (Lemma 4). The proof of existence follows the framework of [Reny,, 1999], which ensures existence by demonstrating that the sum of agents’ payoff functions is upper semi-continuous in $x_{i}\in\mathcal{X}$ and game $G^{\prime}$ is payoff security, which we formally define in the proof. For uniqueness, we build upon the results of [Rosen,, 1965]. Specifically, we first establish the uniqueness of NE in both periods 1 and 2. Then, leveraging the two-agent NE characterization from Lemma 1, we extend these results to the multi-agent setting by iteratively partitioning the agents into two groups and applying the two-agent results. The full proof is provided in the appendix. ∎
This proposition establishes that the quasiconcave CP shaving game $G^{\prime}$ has a unique NE, which always occurs at the agents’ balance points $b_{i}$. Notably, under both concave and quasiconcave conditions, the NE results align with those in the two-agent setting, as the extension from two agents to multiple agents follows a straightforward generalization. This suggests an approach where the NE results of the two-agent non-concave CP shaving game $G$ serve as a foundation for analyzing the $N$-agent non-concave CP shaving game $G^{\prime}$, whose NE is evidently more complex. Fundamentally, Theorem 1 highlights a key insight: system demand is always balanced across the two periods, i.e., $x^{*}+y^{*}=b$. In this process, the less flexible agent—determined by shifting penalty parameters and baseline demand—reaches its critical point first, while the more flexible agent continues adjusting its demand to balance system demand over the two periods.
We define agents whose baseline peak demand coincides with the system baseline CP period as CP-period agents, while those whose baseline peak demand does not align with the system baseline CP period are referred to as non-CP-period agents. The set containing all CP-period agents is denoted as $N_{\mathrm{cp}}$, and the set of all non-CP-period agents as $N_{\mathrm{ncp}}$, such that $N_{\mathrm{cp}}\cup N_{\mathrm{ncp}}=N$ and $N_{\mathrm{cp}}\cap N_{\mathrm{ncp}}=\emptyset$. We now present the following proposition to analyze the equilibrium conditions of the non-concave multi-agent CP shaving game $G^{\prime}$.
@@
NE in the non-concave multi-agent CP shaving game. Given non-concave multi-agent CP shaving game $G^{\prime}$, under the condition of
@@
the hybrid NE solution, where for each CP-period agent and for the whole non-CP-period agent set, is given by
$x_{i}^{*}=\min\{r_{i},b_{i}\},i\in N_{\mathrm{cp}},\sum_{i\in N_{\mathrm{ncp}}}x_{i}^{*}=b-\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}.$ (24b)
Otherwise, the hybrid NE solution, where for each non-CP-period agent and the overall CP-period agent set, is given by
@@
The proof of this proposition is based on introducing two virtual agents, each representing the aggregated baseline demand and shifting penalty conditions of the CP-period and non-CP-period agent sets. While these virtual agents share the same baseline conditions as their respective agent sets, their strategy structure differs, as the agent sets’ strategies result from the aggregation of individual agents’ decisions. Using Theorem 1, we derive the best strategies for the virtual agents. From the best response analysis, we establish that both the agent sets and the virtual agents exhibit the same strategic behavior due to their identical baseline conditions. Specifically, both will shift demand away from or toward the CP period to minimize costs, with cost reduction directly related to shifting amounts. Finally, by linking the virtual agents’ strategies to those of the agent sets, we derive the equilibrium conditions for the CP-period and non-CP-period agent sets. The detailed proof is provided in the appendix. ∎